# Enhancing Fault Localization Through Ordered Code Analysis with LLM Agents and Self-Reflection

### v1

@@ 3. Methodology
Figure 1. An overview of LLM4FL.
Applying Large Language Models (LLMs) for fault localization presents specific challenges, such as token size limitations and performance issues with longer input contexts (Levy et al., 2024; Liu et al., 2024; Wu et al., 2023). Processing large datasets like test coverage and source code in a single query is impractical due to these constraints. Moreover, there are limited studies on effective prompting techniques for fault localization. To address these issues, we propose LLM4FL, an LLM-agent-based fault localization system. Figure 1 illustrates an overview of our technique. LLM4FL operates in four stages: 1) Analyzing the Covered Methods Using a Divide-and-Conquer Strategy, 2) Defining LLM Agent Roles and Tools, 3) Code Navigation through Agent Collaboration and Prompt Chaining, and 4) Self-Reflection and Chain-of-Thought Reasoning. First, the coverage data is divided into smaller, manageable groups to fit within the LLM’s token limits. Second, LLM agents are assigned specific roles and equipped with tools to analyze these groups. Third, LLM agents collaborate through prompt chaining to navigate the code autonomously and identify potential faults. Finally, the agents refine their analysis using self-reflection and chain-of-thought reasoning. Below, we provide more details on each of these stages.
3.1. Analyzing the Covered Methods Using Divide-and-Conquer
Prior studies in fault localization utilized dynamic execution data, such as coverage information from both failing and passing test cases, to identify faulty code locations (Lou et al., 2021; Li et al., 2019; Abreu et al., 2009; Rafi et al., 2024; Qian et al., 2023b). However, Large Language Models (LLMs) face challenges when processing large-scale coverage data due to inherent token size limitations (Kang et al., 2024). These limitations lead to performance degradation, extended processing times, and reduced accuracy when input tokens are truncated. To address these constraints, we propose a divide-and-conquer strategy, dividing the coverage data (i.e., containing the covered methods and the corresponding code) into manageable groups that the LLM can process within its token limits.
We generate a list of covered methods using tools like GZoltar (Campos et al., 2012), focusing only on methods executed by failing tests. To manage the token limitation issue, we include only the specific method statements that were executed by the failing tests, excluding any uncovered portions of the code. Before doing a divide-and-conquer, we use a Spectrum-Based Fault Localization (SBFL) technique to sort the covered methods so that those more likely to be faulty can be grouped together. This approach is inspired by a recent study (Tyen et al., 2024) that demonstrates LLMs perform better when the order of instruction is carefully considered. Note that, regardless of the order, LLM4FL eventually analyzes every covered method. Specifically, we use Ochiai (Abreu et al., 2006) to sort the methods before the divide-and-conquer process. Ochiai is a widely used unsupervised spectrum-based fault localization technique because of its high efficiency and good fault localization accuracy (Abreu et al., 2006; Lou et al., 2021; Li et al., 2021; Cui et al., 2020; Wen et al., 2019; Qian et al., 2021). Intuitively, Ochiai assigns a higher suspiciousness score to a statement that is executed more frequently by failing cases and less frequently by passing test cases. Ochiai is defined as:
$Ochiai(a_{ef},a_{nf},a_{ep})=\frac{a_{ef}}{\sqrt{(a_{ef}+a_{nf})\times(a_{ef}+a_{ep})}},$
where $a_{ef}$ is the number of failed test cases that execute a statement, $a_{nf}$ is the number of failed test cases that do not execute the statement, and $a_{ep}$ is the number of passed test cases that execute the statement. The result of Ochiai is a value between 0 and 1, with higher values indicating a higher likelihood that a particular statement is faulty. We then aggregate the scores for every statement in a method to get the method-level score.
Once the covered methods are sorted, we divide the methods into smaller groups, ensuring that each group fits within the token limitation of the LLM. More formally, let $C$ represent the sorted covered methods, containing a sequence of pairs ($m$, $s$), where $m$ denotes a covered source method and $s$ denotes the corresponding code statements. The group length is determined using the token limitation specified in the official documentation of GPT-4o-mini (the LLM that we use in this study), which has an input context window of 128,000 tokens (OpenAI, 2024). The length is calculated as $TotalTokenLength/TokenLimitation$, ensuring each group remains within the model’s token limitation. Specifically, we divide $C$ into to sequence of $C_{1},C_{2},\ldots,C_{n}$ each $C_{i}$ contains the subset of pairs ($m_{i}$, $s_{i}$), and satisfies the constraint —$C_{i}$$<=TokenLimitation$.
3.2. LLM Agents and the Available Tools
LLM4FL emulates developers’ debugging process by implementing two LLM agents: 1) Tester and 2) Debugger. Below we discuss the details of these agents:
Tester Agent. The tester agent’s main goal is to identify and prioritize suspicious methods. More formally, it navigates through each group $C_{i}$ of the covered methods. The Tester Agent’s input includes the sorted covered methods in each group $C_{i}$ and the corresponding test code, stack traces, and related test information from failing executions. To further reduce the input token size and allow LLM to focus on the relevant code, we only provide the portion of the test code leading to the failure as input. We analyze the stack trace to identify where the test failed and extract the code up to that point, omitting the rest.
Figure 2 gives an example of the Tester Agent’s prompt. The Tester Agent is responsible for identifying suspicious methods likely related to faults in the system. To achieve this, it needs to 1) analyze methods executed by failing tests, 2) review relevant test code, and 3) examine stack trace data to pinpoint where failures occurred. To facilitate this process, we provide the Tester Agent with three tools. The tool get_covered_methods allows the agent to retrieve the source code of the methods executed during failing tests, get_test_code fetches the relevant test code, and get_stacktrace gathers stack trace data. Each tool has a unique name and description, which the agent uses to determine when to apply the tool during its analysis. The agent decides autonomously, based on its task, which tool to invoke by interpreting the tool’s description and relevance to the current stage of fault localization. Additionally, each method within the grouped data is assigned a unique ID during the dividing process, allowing the agent to reference and retrieve specific methods. After analyzing all groups, the Tester Agent generates a list of suspicious methods, denoted as $S_{i}\subseteq M_{i}$, where $M_{i}$ represents the methods in group $C_{i}$. This list of suspicious methods is then consolidated into a final list of method IDs for further analysis.
Prompt Template for Initial Candidate Selection with Advanced Tools As an Intelligent Tester Agent, utilize the following tools to identify methods most likely related to the fault. Focus on analyzing the test code, stack trace, and covered methods to pinpoint suspicious or fault-inducing methods. Ensure the exact method_id is included in the final JSON output. Available Tools:
get_test_code()       – Retrieves the complete test code.
get_stacktrace()         – Access the stack trace of the failed test.
get_covered_methods()   – Lists all methods executed during the test.
Your Task: Use the insights from the tools and provide the potential suspicious methods in the following JSON format:
        "method_ids": [1, 2, 3, ...]  // The potential suspicious method’s ids
Figure 2. Enhanced Prompt for Initial Candidate Selection.
Debugger Agent. Figure 3 gives an example of the Debugger Agent’s prompt. The Debugger Agent is responsible for thoroughly evaluating and ranking a given list of candidate methods $S$. To perform its analysis, the Debugger Agent relies on three tools designed to extract relevant information. get_test_code and get_stacktrace provide LLM’s requested information by retrieving the relevant test code and stack trace data, helping the Debugger Agent understand the failure scenarios associated with each method. Additionally, the get_method_body tool allows the agent to fetch the full source code of a method using its unique method ID, enabling the Debugger Agent to analyze each method’s source code.
Once the Debugger Agent completes its analysis, it assigns a rank score $\sigma_{j}$ to each method $m_{j}$, representing the likelihood that the method is contributing to the fault. The candidate methods in $S$ are then ranked based on these scores, producing a ranked list $R$ that orders the methods from most to least suspicious:
$R=\text{sort}(\{(m_{j},\sigma_{j}):m_{j}\in S\})$
This final ranked list $R$ is output in JSON format for further evaluation and parsing.
Prompt Template for Code Understanding and Fault Analysis As an Advanced Debugger Agent, use the following tools to analyze the provided method bodies based on the Tester Agent’s insights from the test and stack traces. The fault may propagate across multiple methods or be indirectly related to the original issue. Available Tools:
get_test_code()       – Retrieve the entire test code for detailed analysis.
get_stacktrace()         – Access the stack trace to trace method calls and identify potential faults.
get_method_body()      – Retrieve the body of specific methods for further inspection.
Task: Using the provided tools, analyze and focus on the following Methods: {suspicious methods} Rank these methods from most to least suspicious, and provide brief reasoning for each based on its behavior and potential involvement in the fault. IMPORTANT: Provide your output in the following JSON format:
            "method_id": int,  // The most suspicious method’s id
            "reasoning": string,  // Reasoning for its suspiciousness
            "rank": int  // Rank of suspiciousness
Figure 3. Enhanced Prompt for Code Understanding and Fault Analysis.
3.3. Code Navigation Through Prompt Chaining
Code navigation is a critical part of how developers trace faults in real-world scenarios. When debugging, developers often start by analyzing a specific method and then explore its caller or callee methods to better understand the overall logic and pinpoint where the fault might lie. Our approach emulates this process by implementing code navigation through prompt chaining, allowing the Tester and Debugger Agents to collaborate and progressively focus their analysis on the most relevant parts of the codebase.
The process begins with the Tester Agent, which works through each group of the coverage data $C=\{C_{1},C_{2},\ldots,C_{n}\}$, where each group $C_{i}$ contains a subset of the covered methods $M_{i}$. The Tester Agent identifies suspicious methods within each group, producing a list $S_{i}\subseteq M_{i}$ of candidate methods. This step mimics how a developer might flag areas of interest in the code that seem related to the fault. After analyzing all the groups, the Tester Agent compiles the suspicious methods into a consolidated list:
$S=\bigcup_{i=1}^{n}S_{i}$
This list $S$ is passed to the Debugger Agent, initiating the prompt chaining process.
The Debugger Agent takes the output from the Tester Agent and dives deeper into the suspicious methods. In this stage, the Debugger Agent navigates through each method in $S$, inspecting not only the method itself but also the related methods, such as caller and callee methods, if the agent believes further analysis can help identify faulty code location. The Debugger Agent emulates the manual process of tracing how faults propagate through different parts of the code. By leveraging its tools, the Debugger Agent retrieves information on how each method interacts with others and how it may contribute to the fault. If the Debugger Agent identifies that a method $m_{j+1}$ is called within $m_{j}$ and might require further examination, it can request the code for $m_{j+1}$ if it is part of the candidate list, ensuring deeper analysis where necessary. LLM4FL implements an LLM-based fault navigation system that allows the Debugger Agent to dynamically traverse the codebase and retrieve additional relevant methods for further inspection. Through this process, the Debugger Agent traces potential fault propagation pathways and evaluates how faults spread across methods.
3.4. Self-Reflection and Chain-of-Thought for Improved Fault Localization Results
LLM4FL uses two additional prompting techniques to further improve the fault localization results.
Self-Reflection. In our approach, both the Tester and Debugger Agents engage in self-reflection to enhance the quality of their outputs. After completing their initial analysis, each agent operates in two phases: a generator phase and a reflector phase. During the generator phase, the agents produce their primary results, such as identifying suspicious methods or ranking them by suspiciousness. In the reflector phase, the agents critique their own work, offering feedback to refine their results and correct potential errors. While it has not yet been studied in the fault localization context, a recent study (Renze and Guven, 2024) found that self-reflection improves LLM’s performance in problem-solving tasks.
Chain-of-Thought for Refining Fault Localization. We incorporate a chain-of-thought reasoning process to enhance fault localization. After generating a ranked list of suspicious methods, the LLMs propose potential fixes for the top-ranked methods. The agents then re-rank the methods by considering the generated fixes. This process allows the agents to think more deeply about the faults and reassess the ranking based on the insights gained from considering these fixes (Bao et al., 2024; Wang et al., 2024).

### v2

@@ 3. Methodology
Traditionally, developers rely on dynamic execution data to facilitate fault localization: (i) code coverage is used for narrowing down executed code that changes the behavior incorrectly, whereas (ii) stack trace provides developers with insight into the program execution when the exception gets thrown, helping them diagnose the issue more effectively. However, directly applying Large Language Models (LLMs) for fault localization is infeasible. This is due to the large size of software artifacts, which frequently exceed token limitations, leading to input truncation and degraded performance  (Levy et al., 2024; Liu et al., 2024; Wu et al., 2023). To mitigate this issue, we propose LLM4FL, an LLM-agent-based fault localization system, which adopts a divide-and-conquer algorithm, which first divides the large software artifacts, and employs multiple LLM-Agents to localize faults autonomously.
Figure 1 provides an overview of our technique. LLM4FL consists of three LLM-Agents by using novel prompting techniques: (i) Context Extraction Agent, (ii) Debugger Agent, and (iii) Reviewer Agent to localize the fault iteratively. Specifically, the Context Extraction Agent utilizes a order-aware division and failure-reason guided prioritization prompting technique, which consists of two phases. In the division phase, a toolchain divides the large-scale code coverage data into small groups to fit within the LLM’s token limits. In the prioritization phase, the agent iteratively analyzes the divided code coverage, leveraging failed test cases and stack traces to identify potentially faulty methods. The Debugger Agent then performs graph-based retrieval-augmented code navigation prompting to locate code artifacts further to enhance the accuracy of fault ranking. Finally, the Reviewer Agent re-ranks the buggy method through verbal reinforcement learning prompting with the reviewer agent, ensuring more precise fault localization. Below, we provide more details on each of these phases.
Figure 1. Overview of LLM4FL for Multi-Agent Fault Localization, illustrating how agents collaborate to analyze software artifacts, extract failure reasoning, perform graph-based retrieval-augmented code navigation, and rank faulty methods using verbal reinforcement learning.
3.1. Context Extraction Agent
This agent defines a order-aware division and failure-reason guided prioritization prompting technique to divide code-coverage and iteratively prioritize faulty methods using stack trace and test code within each division. We describe the technique below.
3.1.1. Order-Aware Division Phase
To mitigate token size limitation, our LLM4FL first performs a division phase, where it first runs the test case to extract the list of method-level code coverage using GZoltar (Campos et al., 2012), which is denoted as $C$, containing a sequence of pairs ($m$, $s$), where $m$ denotes the method and $s$ denotes the set of statements within method $m$. We then divide $C$ into $K$ group of sequences of $C_{1},C_{2},\ldots,C_{k}$, where $K=\left\lceil\frac{\text{Input Token Length}}{\text{Token Limitation}}\right\rceil$, and each subset satisfies $|C_{i}|\leq\lfloor\text{Token Limitation}\rfloor$ to ensure the data fits within the LLM’s context window. For example, given an Input Token Length of 500K tokens, we divide this by the Token Limitation for GPT-4o-mini, which has a limit of 128K tokens (OpenAI, 2024). This results in $K=4$, each with 128K token limit.
Traditionally, the order in which divisions occur in divide-and-conquer algorithms does not impact the outcome. However, prior studies have shown that LLM may improve performance when the order of instructions is carefully considered  (Chen et al., 2024a). Inspired by this, we propose order-aware division, where we use a Spectrum-Based Fault Localization (SBFL) technique to sort $C$ based on their likelihood of being faulty. Specifically, we use Ochiai ranking to order $C_{1},C_{2},\ldots,C_{k}$, which is an efficient and unsupervised technique that assigns higher suspiciousness scores to statements that are executed more frequently by failing test cases and less frequently by passing ones (Abreu et al., 2006; Lou et al., 2021; Li et al., 2021; Cui et al., 2020; Wen et al., 2019; Qian et al., 2021).
3.1.2. Failure-Reason Guided Prioritization Phase
To prioritize faulty methods from the divided code coverage, one approach is to provide the agent with the stack trace and test case. However, prior studies have shown that incorporating summarized description can improve accuracy (Stiennon et al., 2020; Roit et al., 2023). Thus, we propose failure-reason-guided fault prioritization, which follows two steps. First, the input tokens are summarized into a failure-reason representation, capturing the test purpose, expected output, and failure reason. Second, the divided code coverage is prioritized based on the reason for failure. Below, we discuss each step in detail.
Figure 2. Context-Extraction Agent uses tool-chains to preprocess software artifacts for Lang-5 to emphasize the test failure context. For (i) stack-trace the agent prunes external libraries and (ii) test code, the agent prunes statements in the test code after the assertion failure.
Generate failure-reason from code artifacts. Our failure-reason generation makes the following three observations: (i) Firstly, the stack traces are verbose and include calls to external libraries unrelated to the fault. (ii) Secondly, some statements within a test case are irrelevant to the failure, specifically statements after the first assertion failures. (iii) Finally, the test case may call other helper methods that trigger the test failure (Peng et al., 2022). Hence, as shown in Figure 2, to mitigate (i), the Context Extraction Agent uses PreprocessTrace tool-chain to prune external execution in the stack trace. To mitigate (ii) and (iii), Context Extraction Agent uses PreprocessTest tool-chain, which uses static analysis to build an interprocedural call graph to extract all helper test methods called by the test and prunes all the statements appearing in the test after the failure point.
Given the preprocessed stack trace and test code, we prompt Context-Extraction Agents to generate failure-reason by answering test purpose, expected output, and failure reason. Below, we show a simplified example of the prompt:
“Analyze the test body and stack trace to generate a test failure reason summarizing the issue with expected output, and potential faulty location.”
## Input: ‘
{Test Code, Stack Trace}
## Response: ‘
{Failure-Reason}
An example of failure-reason generated for Lang-5 given stack trace and test code is:
## Test Purpose:
Test whether LocaleUtils.toLocale parses ’_GB’ as a locale string, expecting it to return a Locale object with an empty language and ’GB’ as the country.
## Failure Reason:
The actual output is an IllegalArgumentException due to an invalid locale format, as LocaleUtils.toLocale requires a language before the underscore. For example, ’_GB’ does not conform to the expected locale format. Update the test input to ’en_GB’ or modifying LocaleUtils.toLocale to handle ’_GB’ as a valid case. d
Failure-reason guided prioritization. Given Failure-reason, the Context Extraction Agent analyzes the covered methods in each group, previously denoted as $\{C_{1},C_{2},\dots,C_{K}\}$, to prioritize methods that are most related to the failure. Formally, prioritizing each set $C_{i}$ results in a prioritized subset $C^{\prime}_{i}$, where: $C^{\prime}_{i}\subseteq C_{i},\forall i\in\{1,2,\dots,K\}$. The prioritized methods from all subsets are then combined into a final prioritized set: $C^{\prime}=\bigcup_{i=1}^{K}C^{\prime}_{i}$. This final prioritized list $C^{\prime}$ is just a union of the ordered subsets, meaning no additional processing is introduced. This union process is necessary as LLMs are highly sensitive to initial order (Chen et al., 2024a), yet analyzing entire code coverage poses a token limitation. By dividing coverage into groups, prioritizing each, and unionizing subsets, we extract important context that are necessary for the Debugger Agent for fault analysis and localization in the next step.
3.2. Debugger Agent
Previously, the Context Extraction Agent prioritized $C$ to subsets $C^{\prime}$ to mitigate the LLM’s token limitation when analyzing large coverage data. While this technique can adapt to any LLM and token input size, prioritization may miss methods in the coverage that are important for fault localization. Inspired by how developers analyze the call dependencies to understand program execution and identify faults, we propose a graph-based retrieval-augmented code navigation strategy (Graph-RAG). This approach enables the Debugger Agent to utilize call graphs and method bodies for effective repository navigation, implementation analysis, generation of failure reasoning, and ranking the methods based on their likelihood of causing errors. We will detail this technique below.
3.2.1. Generating failure reasoning through call-graph-aware retrieval-augmented (Graph-RAG) code navigation.
The agent first uses code-coverage $C$ to construct the inter-procedural call-graph database $G=(V,E)$, where $V$ is a set of methods in the call-graph, and $E$ is a set of edges representing the caller-callee relationship. This graph is used by the tool-chains, get_MethodBody and get_CallGraph to facilitate agent-driven navigation within the call graph. LLM4FL then uses the prompt template below to drive the agent:
“Given (i) failure reasoning, (ii) stack trace, and (iii) test code, analyze, navigate, and enhance the failure reasoning for the given methods to identify faults. If additional implementation details are needed, extract the call graph and retrieve the caller or callee’s implementation. At the end, you should output the (i) all of the analyzed method, (ii) enhanced failure reasoning, (iii) fault ranking.”
## Input:
{Test Code, Stack Trace, Failure Reason, PrioritizedMethods}
## Response:
{Analyzed Method, Enhanced Failure Reason, Fault Ranking}
The agent begins the analysis of methods selected sequentially from the prioritized $C^{\prime}$. For each method, the agent retrieves its body using get_MethodBody and analyzes its implementation. It then queries get_CallGraph to identify relevant callers or callees, deciding whether to retrieve their implementations to enhance its understanding of code behavior. Throughout the iterative process of retrieval and analysis, every method examined, regardless of whether it was prioritized or invoked by a prioritized method, is assigned a failure reason by the agent. This information is then saved in a set referred to as $R$. This failure reasoning is a natural language explanation of each method’s failure analyzed in the code navigation process. These results are stored as pairs in a set: $R=\{(m_{i},r_{i})\mid m_{i}\text{ is analyzed},\;i=1,\dots,n\}$. Here, $n$ denotes the number of methods for which failure reasoning was generated.
Ranking methods based on failure reasoning results. Finally, given the failure reasoning $R$, the agent assigns each method $m_{i}\in\mathcal{R}$ an ordinal rank ($r^{*}_{i}$). Formally, the final ranked set $R^{*}$ is sorted in descending order based on the ordinal ranks:
$R^{*}=\text{sort}\left(\{(m_{i},r_{i},r^{*}_{i})\mid m_{i}\in R\}\right)$
where $r*_{i}$ denotes the ordinal rank that reflects each method’s likelihood of containing the fault, and $r_{i}$ explains the reasoning behind the method being faulty. The final ranked list $R^{*}$ is output in JSON format for subsequent analysis.
3.3. Reviewer Agent
Code review is important in software development to ensure software quality. Beyond traditional debugging, code review can help refine failure reasoning and improve fault localization. Inspired by this, we propose Reviewer Agent, which uses a novel prompting technique called Re-ranking through Verbal Reinforcement Learning to emulate a rigorous code-review process to refine fault localization. We describe the technique below.
3.3.1. Re-ranking through Verbal Reinforcement Learning
To refine the fault rankings $R^{*}$, Reviewer Agent adopts a verbal reinforcement learning process inspired by the Reflexion framework (Shinn et al., 2024). The agent iteratively critiques its rank trajectory by identifying inconsistencies, retrieving missing execution insights, and updating method prioritization to better reflect the failure’s root cause. After the iterations, it applies chain-of-thought to update the final ranking and return the fault localization results.
“Given (i) failure reasoning, (ii) stack trace, and (iii) test code, analyze and critique the initial ranking of suspicious methods. Reflect on the reasoning quality and identify missing or superfluous details. If additional insights are needed, extract the call graph and retrieve the caller or callee’s implementation to refine the ranking. At the end, you should output (i) the revised set of analyzed methods, (ii) the improved failure reasoning, and (iii) the final fault ranking with a possible fix.”
## Evaluating the Initial Ranking:
{Test Code, Stack Trace, Ranked List with Reasonings}
## Self-Critique and Refinement:
- Identify missing insights and remove unnecessary information.
- Retrieve method details using ‘get_MethodBody‘ if needed.
- Extract the call graph with ‘get_CallGraph‘ to enhance fault reasoning.
## Final Output:
{Analyzed Methods, Enhanced Failure Reasoning, Fault Ranking}
Evaluating and self-critiquing the initial ranking. The Reviewer Agent uses the initial ranking, $R^{*}$, provided by the Debugger Agent as the baseline trajectory for refinement. The agent enters a self-reflection phase, guided by Reflexion’s Self-Reflection Model ($M_{sr}$), to refine its ranking decisions through verbal reinforcement learning (Shinn et al., 2024). In this phase, the agent examines whether the initial ranking aligns with observed fault rankings by comparing the failure reasoning $r$, with caller-callee interactions and test coverage data, thereby identifying discrepancies such as misordered rankings or missing call dependencies. The agent generates natural language critiques and retrieves additional execution details if needed using tools such as get_MethodBody (to obtain complete method implementations) and get_CallGraph (to retrieve caller-callee relationships). The refined context is then incorporated in the Reviewer Agent’s iterative adjustment process.
Ranking adjustment through trajectory optimization. After the initial self-critique phase, the Reviewer Agent refines its initial ranking by updating each method’s rank $r^{*}$ and revising its failure reasoning $r$. Conceptualized as a trajectory optimization problem within the Reflexion framework, the agent, acting as an Actor ($M_{a}$), leverages reinforcement cues from its Self-Reflection Model ($M_{sr}$) to systematically re-order methods based on the updated evidence of inter-method interactions and failure reasoning. Each iteration integrates new feedback to adjust the ranking until it stabilizes or a preset iteration limit is reached.
Finalizing the ranking through chain-of-thought. After the iterative ranking adjustment, the Reviewer Agent finalizes the ranking one last time using chain-of-thought (Wang et al., 2024). To facilitate the thinking process, we ask the Reviewer Agent to generate a probable fix for every method in the final ranking by considering the updated ranking score $R^{*}$ and refined failure reasoning $R$. After generating all probable fixes, the agent then revisits all the information to do a final re-ranking. Formally, the final ranked list $R^{*}_{final}=\text{sort}\left(\{(m_{i},r^{*}_{i},r_{i},f_{i})\mid m_{i}\in C^{\prime}\}\right)$, where $f_{i}$ denotes the fix generated for method $m_{i}$. The final ranked list $R$, encoded in JSON format, provides a structured, machine-readable prioritization along with human-interpretable failure justifications.